OpenStudy (anonymous):

What is the acceleration of a car moving along a straight line path that increases its speed from zero to 100 km/h in 10s ?

OpenStudy (anonymous):

\[a_{avg}=\frac{ \Delta v }{ \Delta t }\] \[\Delta v=v_f-v_0=100-0=100\frac{ km }{ h}\] \[\Delta t=t_f-t_0=10-0=10s\] Next, we have to convert our units. As we can see, we have two different units of time here: one in hours and one in seconds. So we must convert either one to the other. Either way, we will get the same answer. \[\frac{ 100km }{ hr }*\frac{ 1hr }{ 60\min }*\frac{ 1\min }{ 60s }=0.0278\frac{ km }{ s }\] Now you are ready to plug these values into your equation.

OpenStudy (irishboy123):

 i'm with @CShrx on this one, but note that we need to assume the question requires average acceleration over the journey.